Hello Algebra Tutors,
Can you please help me with the following problems I am struggling with...
1) Which number does not equal 0.86?
- 0.860 or - .086
2) Write one thousand, five hundred six and three tenths as a decimal.
3) Translate in mathematical terms: a loss of 2 yards is better than a loss of 3 yards. (Hint= inequality signs!)
Please email me back as soon as possible for you.
I really need help and I really need to be able to show my work for each problem!
-Thanks.

Answer by KnightOwlTutor(293)

.086 is the answer. This is 86/1000 whereas .860=86/100
The difference between these two numbers is a factor of 10

In order to make these camparable you need to have a common denominator which in this case is 1000 therefore 86/1000 and 860/1000. See the difference?
1506 3/10=1506.3 divide 3/10 to get the appropriate decimal value.

X=some unknown value
x-2>x-3